OpenStudy (anonymous):

Die rolling An ordinary (fair) die is a cube with the numbers through on the sides (represented by painted spots). Imagine that such a die is rolled twice in succession and that the face values of the two rolls are added together. This sum is recorded as the outcome of a single trial of a random experiment. Compute the probability of each of the following events: Event : The sum is greater than 8 . Event : The sum is an odd number

OpenStudy (anonymous):

Total posible outcomes are 36 P(sum greater than 8) = 10/36 = 5/18

OpenStudy (anonymous):

P(sum is an odd number) = 18/36 = 1/2